Solve by the method of variation of parameter y''+y=secx

1
Expert's answer
2022-02-09T17:58:02-0500

The corresponding homogeneous differential equation is


"y''+y=0"

Characteristic (auxiliary) equation


"r^2+1=0"

"r=\pm i"

The general solution of the homogeneous differential equation is


"y_h=c_1 \cos x+c_2\sin x"

"y'=c_1 '\cos x-c_1 \sin x+c_2'\sin x+c_2\cos x"

If

"c_1 '\cos x+c_2'\sin x=0,"

then


"y''=-c_1'\sin x-c_1\cos x+c_2'\cos x-c_2\sin x"

Substitute


"-c_1'\sin x-c_1\cos x+c_2'\cos x-c_2\sin x"

"+c_1 \cos x+c_2\sin x=\dfrac{1}{\cos x}"

We have


"c_1 '\cos x+c_2'\sin x=0,""-c_1'\sin x+c_2'\cos x=\dfrac{1}{\cos x}"



"c_1'=-\dfrac{\sin x}{\cos x}c_2'"

"\dfrac{\sin^2 x}{\cos x}c_2'+c_2'\cos x=\dfrac{1}{\cos x}"



"c_1'=-\dfrac{\sin x}{\cos x}"

"c_2'=1"

Integrate


"c_1=-\int \dfrac{\sin x}{\cos x}dx=\ln (|\cos x|)+C_1"

"c_2=\int dx=x+C_2"

The general solution of the nonhomogeneous differential equation is


"y_h=\ln (|\cos x|) \cos x+x\sin x+C_1\cos x+C_2\sin x"